# QoS-aware State-Augmented Learnable Algorithm for Wireless Coexistence Parameter Management

### v1

@@ Abstract
Index Terms:
5G NR-U, Wi-Fi, coexistence, unlicensed spectrum, quality of service (QoS), medium access delay, fairness.
I Introduction
@@
The rest of the paper is organized as follows: Section II reviews related work on general coexistence strategies, the application of RL to NR-U/Wi-Fi coexistence, and constrained RL techniques. Section III presents the system model and formally defines the CPM problem. In Section IV, we develop a multi-objective RL approach for CPM, which extends our prior work in [12]. Section V introduces a constrained RL formulation, discusses a primal-dual method for solving it, and presents the state-augmented RL framework. This motivates the design of our proposed algorithm, QaSAL-CPM, which is detailed in Section VI along with its deep RL implementation. Section VII applies the CPM framework to the coexistence of 5G NR-U and Wi-Fi in unlicensed spectrum, discusses the performance metrics of interest, the simulation setup, and the numerical results. The paper is concluded in Section VIII.
II Related Works
II-A General Coexistence Studies
@@
II-B Bandit-Based Reinforcement Learning
@@
II-C State-Driven Reinforcement Learning
@@
III System Model and Problem Formulation
III-A System Model
@@
Effective coexistence management between 5G NR-U and Wi-Fi in unlicensed spectrum requires adaptive parameter tuning to balance fair spectrum access and QoS performance (e.g., latency, efficiency, and throughput) for PC1 traffic transmitters. Examples of key coexistence parameters include the Contention Window (CW) size and the Arbitration Inter-Frame Space (AIFS) duration, which directly impact how transmitters of both networks compete for channel access. CPM refers to the dynamic adjustment of these parameters based on real-time network conditions. The objective of CPM is to ensure the QoS requirements for high-priority traffic while maintaining fair and efficient coexistence. Fig. 2 illustrates the parameterized optimization of the CPM policy through a Double Deep Q-Network (DDQN), which will be discussed further in Section VI.
@@ III-B Problem Formulation
IV Multi-Objective Optimization for CPM
@@
V Constrained Reinforcement Learning for CPM
@@
V-A Primal-Dual Approach for Constrained CPM
@@
V-B State-Augmented Constrained Reinforcement Learning
@@
VI Proposed Algorithm: QaSAL-CPM
In light of the state-augmentation approach, we propose the QoS-aware State-Augmented Learnable algorithm for the CPM problem (QaSAL-CPM). Let us consider state $\mathbf{S}_{t}$ at time step $t$ of the $k$-th epoch. Augmentation of the dual variables $\boldsymbol{\lambda}_{k}$ into the state space results in a augmented state $\tilde{\mathbf{S}}_{t}=(\mathbf{S}_{t},\boldsymbol{\lambda}_{k}$). To be distinguished from previously introduced parameterization in Section V, we introduce an alternative parameterization for the CPM policy, in which the CPM decisions are represented via the parameterization $\mathbf{a}(\mathbf{\tilde{S}};\boldsymbol{\tilde{\theta}})$, where $\boldsymbol{\tilde{\theta}}\in\boldsymbol{\tilde{\Theta}}$ denotes the set of parameters of the state-augmented CPM policy. Then, we define the augmented version of the Lagrangian in (7) as
@@ VI Proposed Algorithm: QaSAL-CPM
VI-A Deep Reinforcement Learning for QaSAL-CPM
@@
VII Performance Evaluation and Simulation Results
In this section, we evaluate the proposed CPM algorithms in a realistic coexistence scenario involving 5G NR-U and Wi-Fi systems sharing an unlicensed spectrum. We implemented a Python simulation environment based on SimPy that models the MAC-layer behavior of 5G NR-U and Wi-Fi under saturated traffic conditions [^1] https://github.com/mfasihi/QaSAL_Coexistence_MAC. The simulated wireless environment is inherently dynamic and competitive due to the heterogeneous nature of the technologies involved. While 5G NR-U relies on scheduled transmissions with LBT, Wi-Fi uses a distributed contention-based access mechanism (i.e., CSMA/CA). These differences can lead to severe performance disparities and unpredictable behavior when the medium becomes congested. To investigate the behavior of the proposed methods, we first apply the multi-objective RL based algorithm from Section IV, which we call MORL-CPM, to a representative coexistence setting to demonstrate its limitations in handling competing objectives under hard QoS constraints. Subsequently, we apply the proposed QaSAL-CPM algorithm from Section VI across two distinct scenarios:
@@ VII-A Performance Metrics
VII-A1 Medium Access Delay

### v2

# Mohammad Reza Fasihi, Brian L. Mark Dept. of Electrical and Computer Engineering and Wireless Cyber Center

@@ Abstract
Keywords 5G NR-U  $\cdot$ Wi-Fi  $\cdot$ coexistence  $\cdot$ unlicensed spectrum  $\cdot$ quality of service  $\cdot$ medium access delay  $\cdot$ fairness
1 Introduction
@@
The rest of the paper is organized as follows: Section 2 reviews related work on general coexistence strategies, the application of RL to NR-U/Wi-Fi coexistence, and constrained RL techniques. Section 3 presents the system model and formally defines the CPM problem. In Section 4, we develop a multi-objective RL approach for CPM, which extends our prior work in [12]. Section 5 introduces a constrained RL formulation, discusses a primal-dual method for solving it, and presents the state-augmented RL framework. This motivates the design of our proposed algorithm, QaSAL-CPM, which is detailed in Section 6 along with its deep RL implementation. Section 7 applies the CPM framework to the coexistence of 5G NR-U and Wi-Fi in unlicensed spectrum, discusses the performance metrics of interest, the simulation setup, and the numerical results. The paper is concluded in Section 8.
2 Related Works
2.1 General Coexistence Studies
@@
2.2 Bandit-Based Reinforcement Learning
@@
2.3 State-Driven Reinforcement Learning
@@
3 System Model and Problem Formulation
3.1 System Model
@@
Effective coexistence management between 5G NR-U and Wi-Fi in unlicensed spectrum requires adaptive parameter tuning to balance fair spectrum access and QoS performance (e.g., latency, efficiency, and throughput) for PC1 traffic transmitters. Examples of key coexistence parameters include the Contention Window (CW) size and the Arbitration Inter-Frame Space (AIFS) duration, which directly impact how transmitters of both networks compete for channel access. CPM refers to the dynamic adjustment of these parameters based on real-time network conditions. The objective of CPM is to ensure the QoS requirements for high-priority traffic while maintaining fair and efficient coexistence. Fig. 2 illustrates the parameterized optimization of the CPM policy through a Double Deep Q-Network (DDQN), which will be discussed further in Section 6.
@@ 3.2 Problem Formulation
4 Multi-Objective Optimization for CPM
@@
5 Constrained Reinforcement Learning for CPM
@@
5.1 Primal-Dual Approach for Constrained CPM
@@
5.2 State-Augmented Constrained Reinforcement Learning
@@
6 Proposed Algorithm: QaSAL-CPM
In light of the state-augmentation approach, we propose the QoS-aware State-Augmented Learnable algorithm for the CPM problem (QaSAL-CPM). Let us consider state $\mathbf{S}_{t}$ at time step $t$ of the $k$-th epoch. Augmentation of the dual variables $\boldsymbol{\lambda}_{k}$ into the state space results in a augmented state $\tilde{\mathbf{S}}_{t}=(\mathbf{S}_{t},\boldsymbol{\lambda}_{k}$). To be distinguished from previously introduced parameterization in Section 5, we introduce an alternative parameterization for the CPM policy, in which the CPM decisions are represented via the parameterization $\mathbf{a}(\mathbf{\tilde{S}};\boldsymbol{\tilde{\theta}})$, where $\boldsymbol{\tilde{\theta}}\in\boldsymbol{\tilde{\Theta}}$ denotes the set of parameters of the state-augmented CPM policy. Then, we define the augmented version of the Lagrangian in (7) as
@@ 6 Proposed Algorithm: QaSAL-CPM
6.1 Deep Reinforcement Learning for QaSAL-CPM
@@
7 Performance Evaluation and Simulation Results
In this section, we evaluate the proposed CPM algorithms in a realistic coexistence scenario involving 5G NR-U and Wi-Fi systems sharing an unlicensed spectrum. We implemented a Python simulation environment based on SimPy that models the MAC-layer behavior of 5G NR-U and Wi-Fi under saturated traffic conditions. The simulated wireless environment is inherently dynamic and competitive due to the heterogeneous nature of the technologies involved. While 5G NR-U relies on scheduled transmissions with LBT, Wi-Fi uses a distributed contention-based access mechanism (i.e., CSMA/CA). These differences can lead to severe performance disparities and unpredictable behavior when the medium becomes congested. To investigate the behavior of the proposed methods, we first apply the multi-objective RL based algorithm from Section 4, which we call MORL-CPM, to a representative coexistence setting to demonstrate its limitations in handling competing objectives under hard QoS constraints. Subsequently, we apply the proposed QaSAL-CPM algorithm from Section 6 across two distinct scenarios:
@@ 7.1 Performance Metrics
7.1.1 Medium Access Delay
@@ 7.1.2 Airtime Fairness
The JFI value ranges from $0.5$ to $1$, with values closer to $1$ indicating a more equitable distribution of airtime. This ensures that neither PC1 nor PC3 traffic classes disproportionately dominates spectrum access.
This work initially examines the CPM problem in a general form with multiple objective functions and introduces algorithms to address it. As a case study, the focus is then narrowed to optimizing the JFI between PC1 and PC3 traffic classes, while ensuring that the medium access delay of PC1 transmitter remains below a specified threshold.
Figure 4: Illustration of medium access delay calculation across time steps. Each row represents a different scenario where transmissions (green) and collisions (orange) occur within a given time step.
7.2 System Setting
Next, we detail the implementation setup used for evaluating the proposed CPM framework in the context of 5G NR-U and Wi-Fi coexistence. The network state $\mathbf{S}_{t}$ is designed to comprehensively capture the dynamic environment of 5G NR-U and Wi-Fi coexistence on an unlicensed spectrum. It encodes critical metrics that influence decision-making, including network performance and resource utilization such as the average and smoothed medium access delay of PC1 transmitter, the collision rates, channel utilization airtime ratio, rate of QoS violations, and JFI. Additionally, it tracks trends in delay variation and short-term collision statistics to provide insights into ongoing network conditions.
The CPM decisions $\textbf{a}_{t}=\textbf{a}(\boldsymbol{\textbf{S}_{t}})$ are selected from the set $\{0,1,\ldots,6\}$. The maximum contention window (CW) size is calculated as $\text{CW}_{\text{max, PC}i}=2^{\text{a}_{i,t}+b}-1$, where $\text{a}_{i,t}$ is the CPM decision for priority class $i$, and $b=0$ and $4$ for PC1 and PC3 traffic, respectively. This design directly influences the backoff timing, which affects each transmitter’s access to the channel. Moreover, the objectives are designed as $f_{0}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))=\mathrm{JFI}_{t}$ and $f_{1}(\mathbf{S}_{t},\mathbf{a}(\mathbf{S}_{t}))=\overline{D}_{\text{PC}1,t}$, denoting the JFI among PC1 and PC3 transmitters and the smoothed medium access delay of PC1 transmitter at time $t$, respectively. We define a constraint for medium access delay of PC1 in terms of a threshold $D_{\text{th}}$.
In MORL-CPM, the agent jointly optimizes medium access delay and airtime fairness through a weighted objective function. In contrast, QaSAL-CPM maximizes the airtime fairness between two traffic classes while ensuring that the medium access delay experienced by PC1 transmitter does not exceed a predefined threshold. To learn these policies, we adopt the DDQN with experience replay proposed in Section 6.1. Each training agent receives an augmented state comprising the system observations and, in the case of QaSAL-CPM, the dual variables associated with delay constraints. Learning is conducted in episodes, each representing a fixed-duration coexistence period.
Figure 5: Average medium access delay of gNB PC1, and fairness between PC1 and PC3 transmitters, under the MORL-CPM algorithm across various delay thresholds in Scenario 2. The dashed line represents the QoS threshold $D_{\text{th}}=2~{}\text{ms}.$
7.3 Problem Formulation
We first solve the CPM for coexistence of 5G NR-U and Wi-Fi through the MORL-CPM algorithm. Utilizing the linear scalarization approach, we define the weight vector as $\boldsymbol{w}=\{1-\alpha,\alpha\}$ according to (4). The trade-off parameter $\alpha\in[0,1]$ balances the relative importance of the objectives. Higher values of $\alpha$ put more emphasize on reducing $\overline{D}_{\text{PC}1,t}$, whereas lower values put more weight on JFI. We normalize $\overline{D}_{\text{PC}1,t}$ by $\text{D}_{\text{max}}$ to normalize the delay term to $[0,1]$. This is critical for combining it with JFI, which is inherently bounded in the same range. Therefore, the multi-objective CPM problem in (5) can be formulated as
$\max_{\{\textbf{a}_{t}\}_{t=0}^{T-1}}{\quad\frac{1}{T}\sum_{t=0}^{T-1}{\left\{(1-\alpha)\mathrm{JFI}_{t}+\alpha\left(1-\frac{\overline{D}_{\text{PC}1,t}}{\text{D}_{\text{max}}}\right)\right\}}}.$ (15)
In contrast, the formulation for QaSAL-CPM introduces the parameterized constrained CPM problem as follows:
$\max_{\{\textbf{a}_{t}\}_{t=0}^{T-1}}\quad$ $\frac{1}{T}\sum_{t=0}^{T-1}\mathrm{JFI}_{t},$ (16a)
s.t. $\frac{1}{T}\sum_{t=0}^{T-1}\overline{D}_{\text{PC}1,t}\leq D_{\text{th}}$ (16b)
This constrained formulation is solved via state-augmented RL, where the Lagrange multipliers associated with delay constraints are updated online and included in the agent’s observation space, enabling real-time adaptation to varying traffic dynamics.
7.4 Congestion-Aware Dual Adjustment and Reward Design
To make the agent more responsive to constraint violations without being overly conservative in low-load conditions, we apply a two-part strategy. First, the dual update step size $\eta_{\lambda}$ in (13) is adjusted based on the violation rate, which is the percentage of time steps in which the medium access delay exceeds the threshold:
$\eta_{\lambda,k}=\eta_{\lambda,\text{min}}+\left(\eta_{\lambda,\text{max}}-\eta_{\lambda,\text{min}}\right)\rho_{\text{violation},k},$ (17)
where $\eta_{\lambda,\text{min}}$ and $\eta_{\lambda,\text{max}}$ are minimum and maximum learning rates for dual updates, and $\rho_{\text{violation},k}$ is the ratio of time steps with delay violations over the update window size, $T_{0}$. This makes $\lambda$ update more aggressively when violations are frequent, and less aggressively when the constraint is approximately satisfied. Second, the reward is computed as
$r_{t}=\text{JFI}_{t}-\beta\max\left(0,\frac{D_{\text{th}}-\overline{D}_{\text{PC}1,t}}{D_{\text{th}}}\right),$ (18)
where $\beta$ is a scaling factor that controls the strength of the penalty applied when the delay falls below the threshold, which encourages the agent to utilize the delay budget more effectively in low-congestion scenarios. In our simulations, $\beta$ was set to a value proportional to the delay threshold; in particular, we used $\beta=1,1.5,2$ for $D_{\text{th}}=1$ ms, $2$ ms, and $3$ ms, respectively.
Table 1: Simulation Setup
  Parameter Value
Interaction time $10,000$ episodes
Episode duration $500$ steps
[…]
Discount factor $0.99$
Replay buffer size 100,000
Range of $\epsilon$ 1.0 to 0.01
DQN learning rate $10^{-5}$
Batch size 16
Hidden layers $128\times 128\times 128$
$\lambda_{\text{max}},T_{0},\eta_{\lambda}$ 5.0, 5, [0.01, 0.2]
7.5 Simulation Results and Analysis
This subsection presents the performance of the proposed methods under different traffic and interference scenarios. We first apply MORL-CPM in Scenario 2 and evaluate its performance. Next, we compare QaSAL-CPM with the Primal-Dual approach in Scenario 1. Finally, we analyze the effectiveness of QaSAL-CPM under varying delay thresholds and network congestion levels. The hyperparameters used for simulation are summarized in Table 1. The step duration is selected to be large enough to include several transmission attempts to enable accurate calculation of the medium access delay.
Figure 6: Evolution of smoothed medium access delay for gNB PC1 with Primal-Dual and QaSAL-CPM algorithms in Scenario 1 ($D_{\text{th,PC1}}=2~{}\text{ms}$).
7.5.1 MORL-CPM
Fig. 5 illustrates how the trade-off coefficient $\alpha$ (see eq. 15) in the MORL-CPM framework affects the average medium access delay experienced by a PC1 transmitter and JFI between PC1 and PC3 transmitters from both networks, respectively. As the number of PC3 transmitters increases, the delay performance degrades across all configurations; however, the degree of degradation depends significantly on the choice of $\alpha$. Notably, lower values of $\alpha$, which place greater emphasis on the fairness (as the secondary objective), result in higher delays for the PC1 transmitter—clearly exceeding the delay threshold of 2 ms. Even moderate scalarization settings (e.g., $\alpha=0.5$) fail to consistently enforce the QoS constraint as the network becomes more congested. Only high scalarization values (e.g., $\alpha=0.7$ and $\alpha=0.9$), which strongly prioritize delay minimization for PC1, maintain the delay below or near the required threshold.
These results highlight a key limitation of scalarized multi-objective based optimization in constrained environments: there is no guarantee of strict QoS satisfaction for high-priority traffic, especially under varying network conditions. This limitation motivates the need for more explicit constraint-handling strategies.
[…]
7.5.2 QaSAL-CPM
We compare the performance of the primal-dual approach (i.e., no state augmentation) and the proposed QaSAL method under Scenario 1. Fig. 6 illustrates the evolution of the smoothed medium access delay for gNB PC1 for the Primal-Dual and QaSAL-CPM algorithms, respectively, over 5000 execution episodes, each of duration $1.25$ s. While the Primal-Dual method satisfies the delay constraint on average, it frequently experiences constraint violations, indicating weaker adherence to real-time QoS requirements. The horizontal, dashed red lines in Fig. 6 indicate the $95$-th percentile smoothed delays over the execution duration, which is a commonly used performance metric for URLLC traffic. For the Primal-Dual method, the $95$-th percentile delay well exceeds the $2$ ms delay threshold, whereas it lies below the threshold for QaSAL-CPM.
In contrast, QaSAL-CPM leverages the state-augmentation approach to maintain the delay consistently below the specified threshold which results in significantly better compliance with QoS constraints. The delay dynamics in QaSAL-CPM are directly influenced by the evolution of the dual variable associated with the delay constraint, as defined in equation (16b). Compared to the Primal-Dual approach, QaSAL-CPM demonstrates a more adaptive and responsive adjustment of the dual variable by spiking in response to constraint violations and decaying as the constraint becomes satisfied and enabling more reliable and timely QoS management.
Fig. 7 presents the average medium access delay experienced by the gNB PC1 transmitter, collision rate and airtime efficiency of gNB PC1, and the fairness across PC1 and PC3 transmitters from both networks, under different delay threshold configurations in Scenario 2. The results clearly show that the QaSAL-CPM algorithm effectively maintains the average delay close to the desired delay levels, whereas more relaxed thresholds lead to higher allowable delays, showing the algorithm’s ability to adapt accordingly. The performance under the “No Learning” baseline exhibits increasing delay as the number of contending transmitters grows, validating the importance of a learning-based dynamic adaptation mechanism like QaSAL in managing coexistence. Notably, the algorithm successfully stabilizes delay performance even as the network becomes increasingly congested. The fairness performance under QaSAL-CPM is strongly influenced by the delay threshold configuration. Tighter thresholds, while reducing delay, tend to lead to lower fairness due to the algorithm prioritizing the delay requirement over the airtime distribution. Conversely, more relaxed delay constraints allow for improved fairness, as seen with the $2.0$ ms and $3.0$ ms delay threshold configurations. Note that when the delay threshold is higher, the agent keeps the delay further below it to avoid violations caused by the added congestion that comes from trying to improve fairness. Overall, these results highlight the inherent trade-off between delay and fairness in constrained multi-objective optimization and demonstrate that QaSAL-CPM can flexibly manage these trade-offs through its tunable delay constraints.
In our experiments, we assumed a worst-case scenario in which the PC1 transmitter’s buffer is always full. In a more realistic scenario, there may be multiple PC1 transmitters, each of which transmits intermittently, allowing more airtime for PC3 traffic. In this paper, QaSAL-CPM controls only one parameter, the CW size. More fine-grained QoS control may be achievable by extending QaSAL-CPM to manage both CW and another parameter, such as AIFS number (see [25]). The delay performance could be improved further by adopting an enhanced version of LBT that employs collision resolution, e.g., [26]. We plan to investigate these and related issues in our future work.
8 Conclusions
We investigated the application of reinforcement learning to CPM in heterogeneous wireless networks, with a specific focus on the coexistence of 5G NR-U and Wi-Fi. We proposed a novel state-augmentated, constrained RL solution called QaSAL-CPM that directly embeds dual variables into the state space. This design enables the learning agent to adapt its policy in real time with improved stability and convergence, relative to constrained RL based on conventional primal-dual methods and multi-objective RL.
We validated our approach through extensive simulations under multiple realistic coexistence scenarios. The simulation results demonstrated that QaSAL-CPM not only satisfies the imposed QoS constraints but also provides smoother training dynamics and improved performance consistency compared to the primal-dual method. Overall, the proposed framework provides a flexible and generalizable approach for managing coexistence parameters in dynamic and complex wireless environments, offering a promising direction for future research.